show that y=ln (x) is a solution of xy" + y'=0.


1
Expert's answer
2021-09-01T23:56:09-0400

"y=\ln x" , find

"y'=\frac{1}{x}, y''=-\frac{1}{x^2}"

Substitute in equation

"x\cdot (-\frac{1}{x^2})+\frac{1}{x}=-\frac{1}{x}+\frac{1}{x}=0"

so "y=\ln x" is a solution of "xy''+y'=0".